# Quantum Fisher information as a witness of non-Markovianity and criticality in the spin-boson model

### v2

@@ Abstract
The quantum Fisher information, the quantum analogue of the classical Fisher information, is a central quantity in quantum metrology and quantum sensing due to its connection to parameter estimation and fidelity susceptibility. Using numerically exact methods applied to a paradigmatic open quantum system, the spin-boson model, we calculate both static and dynamical quantum Fisher information matrix elements with respect to spin-bath couplings and magnetic field strengths. As the spin-bath interaction increases, we first show that the coupling-coupling matrix elements relative to the ground state of the Hamiltonian are linked to the entanglement growth and signal the Berezinskii-Kosterlitz-Thouless quantum phase transition through their non-monotonic behavior. Furthermore, we demonstrate that the time-dependent matrix elements can reveal non-Markovian effects as well as the transition from the coherent to incoherent regime at the Toulouse point, remaining robust under pure dephasing noise. Non-monotonic signatures of the quantum Fisher information matrix reflect changes in quantum resources such as entanglement and coherence, quantify non-Markovian behavior, and enable criticality-enhanced quantum sensing, thereby shedding light on key features of open quantum systems.
@@
In addition to parameter estimation, the QFIM has deep significance in understanding the geometric structure of the Hilbert space, as it quantifies the distinguishability between quantum states corresponding to different parameter values. The Bures distance can be defined as $D_{B}^{2}(\rho_{1},\rho_{2})=2(1-f(\rho_{1},\rho_{2}))\;,$ where $f(\rho_{1},\rho_{2})$ is Uhlmann’s fidelity [12] between the quantum states $\rho_{1}$ and $\rho_{2}$. Considering the distance between infinitesimally close states, one can define the Bures metric matrix $g_{\mu\nu}$. It is found that the Bures metric is proportional to the QFIM: $4g_{\mu\nu}=F_{\mu\nu}\;$ [2, 13]. This relation establishes the QFIM as the metric tensor in the parameter space of quantum states, quantifying state distinguishability and sensitivity under parameter variations. This property makes it central for quantum sensing, as it indicates how to achieve maximum parameter sensitivity, which is relevant for the development of quantum detectors [14, 15, 16, 17].
Near a quantum phase transition (QPT) [18], the ground state is maximally sensitive to the variation of its driving parameter. Indeed, the QFI can also be employed to characterize QPTs and criticality-enhanced sensing [19, 20, 21, 22, 23]. Using the definition of the Bures distance, a strong connection is found between fidelity susceptibility and the QFIM:
@@
In this work, through numerically exact methods, we analyze the QFIM of a paradigmatic open quantum system, the spin-boson model (SBM) [34, 26, 27], which describes the interaction of a spin one-half, representing a generic two-level system (TLS), with a bosonic bath of $N$ harmonic oscillators. This model, which includes amplitude damping processes, exhibits a Berezinskii-Kosterlitz-Thouless (BKT) QPT for strong spin-bath interaction [35, 36, 37]. First of all, we show that the maximum of the QFIM of the Hamiltonian ground state (GS) turns out to be a marker of the related QPT, signaling criticality-enhanced sensing to very weak magnetic fields. Moving from the weak to intermediate coupling regime, the coupling-coupling QFIM element shows a decreasing monotonic behavior linked to increasing von Neumann entropy. Furthermore, we demonstrate that the QFIM dynamics witness the coherent-incoherent transition and provide useful insights into the non-Markovian character of the system through the QFI flow. In particular, in the coherent regime, this flow exhibits characteristic oscillations, whose frequencies are different from those of spin observables.
@@
Model and Numerical Methods.  We focus on the zero-temperature regime and adopt units where $\hbar=1$. The Hamiltonian of the generalized SBM is written as [34]:
@@
where: $1)$ $H_{Q}=-\frac{\Delta}{2}\sigma_{x}-\frac{h}{2}\sigma_{z}$ is the spin one-half free Hamiltonian, with $\Delta$ (our energy unit) and $h$ (typically small) representing magnetic fields along the $x$ and $z$ axes, respectively; $2)$ $H_{B}=\sum_{i=1}^{N}\omega_{i}\bigg{(}b_{i}^{\dagger}b_{i}+\frac{1}{2}\bigg{)}$ is the bath Hamiltonian, a set of $N$ harmonic oscillators with frequencies $\omega_{i}$; $3)$ $H_{I}=\sum_{i=1}^{N}(\sigma_{z}\lambda_{i}^{z}+\sigma_{x}\lambda_{i}^{x})(b_{i}+b_{i}^{\dagger})$ is the qubit-bath interaction Hamiltonian, with couplings $\lambda_{i}^{x}$ and $\lambda_{i}^{z}$. Here, $\sigma_{z}$ and $\sigma_{x}$ are the Pauli matrices representing the TLS, and $b_{i}$ and $b_{i}^{\dagger}$ the annihilation and creation operators for bosonic modes of frequency $\omega_{i}$. The couplings $\lambda_{i}^{z}$ are obtained by discretizing a continuous spectral density function $J(\omega)=\sum_{i}^{N}|\lambda_{i}^{z}|^{2}\delta(\omega-\omega_{i})=\frac{\alpha}{2}\omega_{c}^{1-s}\omega^{s}\Theta(\omega-\omega_{c})$, where $\omega_{c}$ is the cutoff frequency, that we set to $\omega_{c}=10\Delta$, $\alpha$ is the dimensionless spin-bath coupling associated with amplitude damping processes, $s$ is the Ohmicity parameter which we set to $s=1$ (Ohmic bath). The couplings $\lambda_{i}^{x}$ are sampled analogously, and we denote the dimensionless strength of pure dephasing processes by $\kappa$.
We will compute the Hamiltonian GS and dynamics of the QFIM elements by using analytical approximations as well as numerical techniques. Analytical treatments of the dynamics considered in this work rely on Lindblad’s quantum master equation [38] (within this approximation, the static stationary QFIM elements are erroneously zero [39]). This approach is valid at weak couplings for Markovian dynamics, but it breaks down when the system-bath interaction strengthens and non-Markovian effects emerge [26, 27].
@@
Static QFIM through DMRG.  We first focus on the QFIM calculated for the Hamiltonian GS in the absence of pure dephasing term ($\kappa=0$). We analyze the QFIM element $F_{\alpha\alpha}$ (see panel a of Fig. 1), $F_{\alpha\Delta}$, and $F_{\Delta\Delta}$ (see panels a and b of Fig. (1) in the Supplemental Material [39]) as functions of the spin-bath coupling $\alpha$ for small positive values of the magnetic field $h$. The limit $h\rightarrow 0^{+}$ is relevant due to the symmetry breaking that occurs in these systems. Actually, the SBM is known to exhibit a QPT [43, 35, 25, 44, 45] from a delocalized quantum to a localized classical spin state at the critical value $\alpha_{c}\approx 1+O(\Delta/\omega_{c})$ ($\alpha_{c}\approx 1.05$ for our choice of parameters).
@@
For any small $h$, all three QFIM elements display a pronounced peak at strong coupling, meaning that the system becomes increasingly sensitive to variations in the coupling $\alpha$. With decreasing $h$, near criticality, the peak of the QFI $F_{\alpha\alpha}$ shifts towards larger values of $\alpha$ and becomes sharper and higher. Hence, for $h=0^{+}$, $F_{\alpha\alpha}$ is expected to diverge at the exact value $\alpha=\alpha_{c}$. As discussed in the Supplemental Material [39], the same behavior applies to the other two QFIM elements $F_{\alpha\Delta}$ and $F_{\Delta\Delta}$. Indeed, as the spin transitions from the quantum delocalized to the classical localized state, GSs corresponding to arbitrarily close values of $\alpha$ on opposite sides of the transition become nearly orthogonal, leading to a divergence of the QFIM. In the Supplemental Material [39], we detail the procedure used to extract the critical value $\alpha_{c}$ directly from QFIM elements and based on the position of their maxima. With this method we obtain a critical coupling $\alpha_{c}=1.03\pm 0.03$, in excellent agreement with values reported in the literature [35] (the extrapolations from $F_{\Delta\Delta}$ yield similar results).
@@
Up to this point, we have restricted our analysis to the case $\kappa=0$ (zero pure dephasing term). We now examine the QFIM element $F_{\alpha\alpha}$ of the GS of the full Hamiltonian in Eq. (4). As illustrated in Fig. 2, under weak dephasing, the system retains high sensitivity in the vicinity of the critical point for strong $\alpha$. One can interpret these results making a comparison with the plots of panel a of Fig. 1, considering that the spin-bath coupling $\kappa$ in the Hamiltonian of Eq. (4) effectively acts as an additional dynamical external field along $\sigma_{z}$, hence reducing the height of the maximum as $\kappa$ increases. These results are further supported by the behavior of $\langle\sigma_{z}\rangle$ reported in Fig. 6 of the End Matter for the same model parameters and are similar to those obtained for QFIM elements $F_{\Delta\Delta}$, $F_{\alpha\Delta}$ [39]. Importantly, $F_{\alpha\alpha}$ continues to diverge as $1/\alpha$ in the regime of very weak spin-bath coupling $\alpha$, demonstrating that the non-monotonic profile of the QFI as a function of $\alpha$ persists even with small but finite dephasing ($\kappa>0$).
@@
Dynamical QFIM.
@@
We now turn to the analysis of the information encoded in the time evolution of the QFIM. In particular, we focus on the evolution of coupling-coupling element $F_{\alpha\alpha}(t)$, computed using both the numerically exact MPS approach [47] and the Markovian Lindblad master equation, the latter being valid only in the weak spin-bath coupling regime [48]. Within the framework of Lindblad’s approximation, the analytical form of the density operator $\rho(t)$ allows us to derive a closed-form expression for the QFIM [39]. In Fig. 3, we show the QFI element $F_{\alpha\alpha}$ as a function of time for the spin-bath amplitede couplings $\alpha=0.1,0.25,0.5$ and dephasing coupling $\kappa=0$ in the case of $h=0$ (in this regime of parameters, very small values of $h$ introduce only negligible corrections). In the End Matter, we analyze also the role played by finite values of dephasing coupling $\kappa>0$.
@@
The dynamics of the SBM exhibit a coherent-incoherent transition at the Toulouse point $\alpha=0.5$: for $\alpha<0.5$, the system displays coherent oscillations, whereas for $\alpha>0.5$ the behavior becomes incoherent [49]. This can be seen in the dynamics of $\langle\sigma_{z}\rangle$ that shows these two distinct behaviors [45], as shown in Fig. 7 of the End Matter for the same parameters as in Fig. 3.
@@
We first examine the coherent regime (see panels a and b of Fig. 3). For $\alpha=0.1$, MPS and Lindblad predictions of QFI $F_{\alpha\alpha}$ are comparable only at short times, differing significantly as the stationary state is approached. Additional results reported in the End Matter show that even for $\alpha=0.01$, where the two approaches nearly coincide initially, they still deviate at long times (see Fig. 5). Actually, the behavior of $F_{\alpha\alpha}(t)$ within Lindblad approximation can be separated into two regimes: an initial transient characterized by quadratic growth, $F_{\alpha\alpha}(t)\sim t^{2}$, followed by a rapid exponential decay towards zero [48]. Indeed, the Lindblad QFI always vanishes at long times, since the corresponding steady state (the Gibbs zero temperature state) is independent of $\alpha$. In contrast, the MPS QFI dynamics exhibit pronounced oscillations modulated by an exponentially decaying envelope. These oscillations progressively damp out, and the QFI converges to a finite stationary value, in excellent agreement with the DMRG GS results and, consequently, strongly dependent on $\alpha$. As the coupling strength $\alpha$ increases, both the frequency of oscillations and the peak amplitude of the QFI decrease, indicating a reduction in the system’s sensitivity. In this coherent regime, the QFIM flow [33, 1], defined as $\partial_{t}F_{\mu,\nu}$, can be used to characterize non-Markovian effects after the initial transient. As detailed in the End Matter, starting from the maximum of the respective curves, the MPS solution exhibits alternating positive and negative flow, evidencing continuous information backflow between system and bath (see Fig. 8 of the End Matter). By contrast, the Lindblad flow remains strictly negative, indicating a purely outward flow of information.
@@
In the coherent regime, the QFI dynamics are correlated with those of the spin observables. From the comparison with results of $\langle\sigma_{z}\rangle(t)$ reported in the End Matter in Fig. 7, we find that each local maximum of $F_{\alpha\alpha}$ corresponds approximately to an extremum of $\langle\sigma_{z}\rangle(t)$. These points represent regions of high sensitivity to variations in $\alpha$, and therefore coincide with the peaks of the QFI. This interpretation is further supported by the Fourier analysis presented in Fig. 8 of the End Matter: the dominant frequency component of $F_{\alpha\alpha}(t)$ is exactly twice that of $\langle\sigma_{z}\rangle(t)$, a property which remains robust upon pure dephasing noise.
@@
Conclusions.  We have analyzed static and dynamical QFIM in a generalized SBM, a paradigmatic prototype of open quantum systems. When turning on the spin-bath coupling up to the intermediate regime, the QFIM element $F_{\alpha\alpha}$ exhibits a divergence followed by a characteristic monotonic decrease, correlated with an increase of the von Neumann entropy. In the strong coupling regime, the static QFIM reveals clear signatures of criticality-enhanced sensing near the BKT QPT, allowing to identify the values of the critical coupling $\alpha_{c}$. In the coherent regime, the dynamical QFIM displays a characteristic oscillatory behavior whose dominant frequency is exactly twice that of spin observables, establishing a direct connection between QFI flow and non-Markovian effects. Moreover, the dynamical QFIM reflects the coherent-to-incoherent transition that occurs as the spin interaction with the environment increases. Static and dynamic findings remain robust under pure dephasing noise proving that non-monotonic signatures of the QFIM provide powerful probes for identifying criticality, quantifying non-Markovian behavior, and understanding key phenomena in open quantum systems.

### v3

@@ Abstract
The quantum Fisher information, the quantum analogue of the classical Fisher information, is a central quantity in quantum metrology and quantum sensing due to its connection to parameter estimation and fidelity susceptibility. Using numerically exact methods applied to a paradigmatic open quantum system, the spin-boson model, we calculate both static and dynamical quantum Fisher information matrix elements with respect to spin-bath couplings and magnetic field strengths. As the spin-bath interaction increases, we first show that the coupling-coupling matrix elements relative to the ground state of the Hamiltonian are linked to the entanglement growth and signal the Berezinskii-Kosterlitz-Thouless quantum phase transition through their non-monotonic behavior. We also point out that the static quantum Fisher information exhibits a non-perturbative behavior in the zero-coupling limit, which we justify with an analytic argument. Furthermore, we demonstrate that the time-dependent matrix elements can reveal non-Markovian effects as well as the transition from the coherent to incoherent regime at the Toulouse point, remaining robust under pure dephasing noise. Non-monotonic signatures of the quantum Fisher information matrix reflect changes in quantum resources such as entanglement and coherence, quantify non-Markovian behavior, and enable criticality-enhanced quantum sensing, thereby shedding light on key features of open quantum systems.
@@ I Introduction
In addition to parameter estimation, the QFIM has deep significance in understanding the geometric structure of the Hilbert space, as it quantifies the distinguishability between quantum states corresponding to different parameter values. The Bures distance can be defined as $D_{B}^{2}(\rho_{1},\rho_{2})=2(1-f(\rho_{1},\rho_{2}))\;,$ where $f(\rho_{1},\rho_{2})$ is Uhlmann’s fidelity [12] between the quantum states $\rho_{1}$ and $\rho_{2}$. Considering the distance between infinitesimally close states, one can define the Bures metric matrix $g_{\mu\nu}$. It is found that the Bures metric is proportional to the QFIM: $4g_{\mu\nu}=F_{\mu\nu}\;$ [2, 13]. This relation establishes the QFIM as the metric tensor in the parameter space of quantum states, quantifying state distinguishability and sensitivity under parameter variations. This property makes it central for quantum sensing, as it indicates how to achieve maximum parameter sensitivity, which is relevant for the development of quantum detectors [14, 15, 16, 17]
Near a quantum phase transition (QPT) [18], the ground state is maximally sensitive to the variation of its driving parameter. Indeed, the QFI can also be employed to characterize QPTs along with other quantum-information-related quantities, such as entanglement and fidelity [19, 20, 21, 22], and criticality-enhanced sensing [23, 24, 25, 26, 27]. Although the QFI can be related to other quantities, it provides a complementary, explicitly metrological perspective in affording direct access to state sensitivity. Using the definition of the Bures distance, a strong connection is found between fidelity susceptibility and the QFIM:
@@
In this work, through numerically exact methods, we analyze the QFIM of a paradigmatic open quantum system, the spin-boson model (SBM) [37, 29, 30], which describes the interaction of a spin one-half, representing a generic two-level system (TLS), with a bosonic bath of $N$ harmonic oscillators. This model, which includes amplitude damping processes, exhibits a Berezinskii-Kosterlitz-Thouless (BKT) QPT for strong spin-bath interaction [38, 39, 40]. First of all, we show that the maximum of the QFIM of the Hamiltonian ground state (GS) turns out to be a marker of the related QPT, signaling criticality-enhanced sensing to very weak magnetic fields. Moving from the weak to intermediate coupling regime, the coupling-coupling QFIM element shows a decreasing monotonic behavior linked to increasing von Neumann entropy. We uncover a non-perturbative divergence of the QFI in the zero-coupling limit and validate it by benchmarking our numerical data against a perturbative analytic expansion. Furthermore, we demonstrate that the QFIM dynamics witness the coherent-incoherent transition and provide useful insights into the non-Markovian character of the system through the QFI flow. In particular, in the coherent regime, this flow exhibits characteristic oscillations, whose frequencies are different from those of spin observables.
@@
The paper is structured as follows. In Sec. II we present the spin-boson model analyzed in this work, along with the analytical and numerical techniques employed. Our results are reported in Sec. III, where we analyze both the static QFIM in Subsec. III.1 and its dynamical behavior in Subsec. III.2. Finally, we conclude and discuss our findings, together with possible perspectives, in Sec. IV. In the appendices, we provide additional details on , on the static behavior of different QFIM elements and the estimation of the critical coupling (App. A), weak and zero coupling to the bath (App. B), on the static and dynamical features of the $z$ component of the Bloch vector (App. C), on the dynamical QFIM and its relation to non-Markovianity (App. D) and on the Lindblad approximation (App. E).
II Model and Methods
We focus on the zero-temperature regime and adopt units where $\hbar=1$. The Hamiltonian of the generalized SBM is written as [37]:
@@ II Model and Methods
where: $1)$ $H_{Q}=-\frac{\Delta}{2}\sigma_{x}-\frac{h}{2}\sigma_{z}$ is the spin one-half free Hamiltonian, with $\Delta$ (our energy unit) and $h$ (typically small) representing magnetic fields along the $x$ and $z$ axes, respectively; $2)$ $H_{B}=\sum_{i=1}^{N}\omega_{i}\bigg(b_{i}^{\dagger}b_{i}+\frac{1}{2}\bigg)$ is the bath Hamiltonian, a set of $N$ harmonic oscillators with frequencies $\omega_{i}$; $3)$ $H_{I}=\sum_{i=1}^{N}(\sigma_{z}\lambda_{i}^{z}+\sigma_{x}\lambda_{i}^{x})(b_{i}+b_{i}^{\dagger})$ is the qubit-bath interaction Hamiltonian, with couplings $\lambda_{i}^{x}$ and $\lambda_{i}^{z}$. Here, $\sigma_{z}$ and $\sigma_{x}$ are the Pauli matrices representing the TLS, and $b_{i}$ and $b_{i}^{\dagger}$ the annihilation and creation operators for bosonic modes of frequency $\omega_{i}$. The couplings $\lambda_{i}^{z}$ are obtained by discretizing a continuous spectral density function $J(\omega)=\sum_{i}^{N}|\lambda_{i}^{z}|^{2}\delta(\omega-\omega_{i})=\frac{\alpha}{2}\omega_{c}^{1-s}\omega^{s}\Theta(\omega_{c}-\omega)$, where $\omega_{c}$ is the cutoff frequency, that we set to $\omega_{c}=10\Delta$, $\alpha$ is the dimensionless spin-bath coupling associated with amplitude damping processes, $s$ is the Ohmicity parameter which we set to $s=1$ (Ohmic bath). The couplings $\lambda_{i}^{x}$ are sampled analogously, and we denote the dimensionless strength of pure dephasing processes by $\kappa$.
We will compute the Hamiltonian GS and dynamics of the QFIM elements by using analytical approximations as well as numerical techniques. Analytical treatments of the dynamics considered in this work rely on Lindblad’s quantum master equation [41] (within this approximation, the static stationary QFIM elements are erroneously zero App. E). This approach is valid at weak couplings for Markovian dynamics, but it breaks down when the system-bath interaction strengthens and non-Markovian effects emerge [29, 30].
@@
III Results
In the following, we present the results obtained by computing QFIM at equilibrium (DMRG) and out of equilibrium (time-dependent MPS simulations and the Lindblad equation).
III.1 Static QFIM through DMRG
We first focus on the QFIM calculated for the Hamiltonian GS in the absence of pure dephasing term ($\kappa=0$). We analyze the QFIM element $F_{\alpha\alpha}$ (see panel a of Fig. 1), $F_{\alpha\Delta}$, and $F_{\Delta\Delta}$ (see panels a and b of Fig. 4 in App. A) as functions of the spin-bath coupling $\alpha$ for small positive values of the magnetic field $h$. The limit $h\rightarrow 0^{+}$ is relevant due to the symmetry breaking that occurs in these systems. Actually, the SBM is known to exhibit a QPT [45, 38, 28, 46, 47] from a delocalized quantum to a localized classical spin state at the critical value $\alpha_{c}\approx 1+O(\Delta/\omega_{c})$ ($\alpha_{c}\approx 1.05$ for our choice of parameters).
@@ III.1 Static QFIM through DMRG
For any small $h$, all three QFIM elements display a pronounced peak at strong coupling, meaning that the system becomes increasingly sensitive to variations in the coupling $\alpha$. With decreasing $h$, near criticality, the peak of the QFI $F_{\alpha\alpha}$ shifts towards larger values of $\alpha$ and becomes sharper and higher. Hence, for $h=0^{+}$, $F_{\alpha\alpha}$ is expected to diverge at the exact value $\alpha=\alpha_{c}$. As discussed in App. A, the same behavior applies to the other two QFIM elements $F_{\alpha\Delta}$ and $F_{\Delta\Delta}$. Indeed, as the spin transitions from the quantum delocalized to the classical localized state, GSs corresponding to arbitrarily close values of $\alpha$ on opposite sides of the transition become nearly orthogonal, leading to a divergence of the QFIM. In App. A, we detail the procedure used to extract the critical value $\alpha_{c}$ directly from QFIM elements and based on the position of their maxima. With this method we obtain a critical coupling $\alpha_{c}=1.03\pm 0.03$, in excellent agreement with values reported in the literature [38] (the extrapolations from $F_{\Delta\Delta}$ yield similar results).
@@
Up to this point, we have restricted our analysis to the case $\kappa=0$ (zero pure dephasing term). We now examine the QFIM element $F_{\alpha\alpha}$ of the GS of the full Hamiltonian in Eq. (4). As illustrated in Fig. 2, under weak dephasing, the system retains high sensitivity in the vicinity of the critical point for strong $\alpha$. One can interpret these results making a comparison with the plots of panel a of Fig. 1, considering that the spin-bath coupling $\kappa$ in the Hamiltonian of Eq. (4) effectively acts as an additional dynamical external field along $\sigma_{z}$, hence reducing the height of the maximum as $\kappa$ increases. These results are further supported by the behavior of $\langle\sigma_{z}\rangle$ reported in Fig. 8 of the App. C for the same model parameters and are similar to those obtained for QFIM elements $F_{\Delta\Delta}$, $F_{\alpha\Delta}$ (see App. A). Importantly, $F_{\alpha\alpha}$ continues to diverge as $1/\alpha$ in the regime of very weak spin-bath coupling $\alpha$, demonstrating that the non-monotonic profile of the QFI as a function of $\alpha$ persists even with small but finite dephasing ($\kappa>0$).
@@
III.2 Dynamical QFIM and its flux to witness non-Markovianity
@@
We now turn to the analysis of the information encoded in the time evolution of the QFIM. In particular, we focus on the evolution of coupling-coupling element $F_{\alpha\alpha}(t)$, computed using both the numerically exact MPS approach [49] and the Markovian Lindblad master equation, the latter being valid only in the weak spin-bath coupling regime [50]. Within the framework of Lindblad’s approximation, the analytical form of the density operator $\rho(t)$ allows us to derive a closed-form expression for the QFIM, as detailed in App. E. In Fig. 3, we show the QFI element $F_{\alpha\alpha}$ as a function of time for the spin-bath amplitude couplings $\alpha=0.1,0.25,0.5$ and dephasing coupling $\kappa=0$ in the case of $h=0$ (in this regime of parameters, very small values of $h$ introduce only negligible corrections). In the App. D, we analyze also the role played by finite values of dephasing coupling $\kappa>0$.
@@
The dynamics of the SBM exhibit a coherent-incoherent transition at the Toulouse point $\alpha=0.5$: for $\alpha<0.5$, the system displays coherent oscillations, whereas for $\alpha>0.5$ the behavior becomes incoherent [51]. This can be seen in the dynamics of $\langle\sigma_{z}\rangle$ that shows these two distinct behaviors [47], as shown in Fig. 9 of the App. C for the same parameters as in Fig. 3.
@@
We first examine the coherent regime (see panels a and b of Fig. 3). For $\alpha=0.1$, MPS and Lindblad predictions of QFI $F_{\alpha\alpha}$ are comparable only at short times, differing significantly as the stationary state is approached. Additional results reported in the App. B show that even for $\alpha=0.01$, where the two approaches nearly coincide initially, they still deviate at long times (see Fig. 7). Actually, the behavior of $F_{\alpha\alpha}(t)$ within Lindblad approximation can be separated into two regimes: an initial transient characterized by quadratic growth, $F_{\alpha\alpha}(t)\sim t^{2}$, followed by a rapid exponential decay towards zero [50]. Indeed, the Lindblad QFI always vanishes at long times, since the corresponding steady state (the Gibbs zero temperature state) is independent of $\alpha$. In contrast, the MPS QFI dynamics exhibit pronounced oscillations modulated by an exponentially decaying envelope. These oscillations progressively damp out, and the QFI converges to a finite stationary value, in excellent agreement with the DMRG GS results and, consequently, strongly dependent on $\alpha$. As the coupling strength $\alpha$ increases, both the frequency of oscillations and the peak amplitude of the QFI decrease, indicating a reduction in the system’s sensitivity. In this coherent regime, the QFIM flow [36, 1], defined as $\partial_{t}F_{\mu,\nu}$, can be used to characterize non-Markovian effects after the initial transient. As detailed in App. D, starting from the maximum of the respective curves, the MPS solution exhibits alternating positive and negative flow, evidencing continuous information backflow between system and bath (see Fig. 10). By contrast, the Lindblad flow remains strictly negative, indicating a purely outward flow of information.
@@
In the coherent regime, the QFI dynamics are correlated with those of the spin observables. From the comparison with results of $\langle\sigma_{z}\rangle(t)$ reported in App. C in Fig. 9, we find that each local maximum of $F_{\alpha\alpha}$ corresponds approximately to an extremum of $\langle\sigma_{z}\rangle(t)$. These points represent regions of high sensitivity to variations in $\alpha$, and therefore coincide with the peaks of the QFI. This interpretation is further supported by the Fourier analysis presented in Fig. 10 of the App. D: the dominant frequency component of $F_{\alpha\alpha}(t)$ is exactly twice that of $\langle\sigma_{z}\rangle(t)$, a property which remains robust upon pure dephasing noise.
@@
IV Conclusions
We have analyzed static and dynamical QFIM in a generalized SBM, a paradigmatic prototype of open quantum systems. The QFIM element $F_{\alpha\alpha}$ exhibits, in the limit of zero spin-bath coupling, a divergent behavior, verified by analytic calculations, while, into the intermediate coupling regime, it is characterized by a monotonic decrease, correlated with an increase in the von Neumann entropy.
In the strong coupling regime, the static QFIM reveals clear signatures of criticality-enhanced sensing near the BKT QPT, allowing to identify the values of the critical coupling $\alpha_{c}$.
In the coherent regime, the dynamical QFIM displays a characteristic oscillatory behavior whose dominant frequency is exactly twice that of spin observables, establishing a direct connection between QFI flow and non-Markovian effects. Moreover, the dynamical QFIM reflects the coherent-to-incoherent transition that occurs as the spin interaction with the environment increases. Static and dynamic findings remain robust under pure dephasing noise proving that non-monotonic signatures of the QFIM provide powerful probes for identifying criticality, quantifying non-Markovian behavior, and understanding key phenomena in open quantum systems.